# Low-dimensional tori in Calogero–Moser–Sutherland systems

### v1

@@ 2.3.1. The generic stratum
where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. with $\mathrm{det}\,X=1$ according to
The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
@@
(34) $\omega=\exp\left(\frac{2\pi i}{n}k\right)\cdot 1_{n},\qquad\text{ where }k=0,\ldots,n-1.$
@@
If $\gamma_{j}$ are eigenvalues of the matrix $g$, such multiplication changes $\gamma_{j}\mapsto\omega\gamma_{j}$. Eigenvalues of $g$ are exponents of coordinates of Calogero–Moser–Sutherland model $\gamma_{j}=\exp(iq_{j})$. Thus the multiplication of $g$ by $\omega$ shifts coordinates as
@@
Thus, we described coordinates on the generic (maximal) stratum: real coordinates $\{x_{i}\}$ and $\{\mathrm{arg}\,\alpha_{i}\}$ with satisfying constrains $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\prod_{i=1}^{n}\alpha_{i}=1$, $|\alpha_{i}|=1$.
Now let us describe all strata and let us show that $(x,\alpha)$ are the action-angle variables.
Define $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
@@
Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\log\alpha_{k}$ for $j=2,\ldots,n$. Then,
@@
Thus, we get coordinates on the generic (maximal) stratum: real coordinates $(y_{2}\ldots,y_{n})$ satisfying constrains $y_{i}>c$ and $(\theta_{2}\ldots,\theta_{n})$. In section 4 we show that $(y,\theta)$ are the action-angle variables.
3. Low–dimensional tori
The moment map equation (6) when $x$ is diagonal becomes
(38) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
where $\lambda=g\bar{\psi}$.
Proposition 1.
If for some $k$ we have $\psi_{k}=0$, then in the $k$-th column and $(k-1)$-th row of the martrix $g$ only nonzero element is $g_{k-1,k}\neq 0$ and $|g_{k-1,k}|=1$.
Proof.
Consider equation (38) for $j=k$
(39) $(x_{k}-x_{i}+c)g_{ik}=0.$
If for all $i$ elements $g_{ik}=0$ then $g$ not invertible. Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$. For $j\neq k-1$ $x_{j}-x_{k}+c\neq 0$ and therefore $g_{jk}=0$. This means that $k$-th column has only one non-zero element $g_{k-1,k}$.
Unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the $(k-1)$-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$. ∎
Assume that $\psi_{j}\neq 0$ only for $j=1,j_{1},\dots,j_{s}$ where $j_{1}<j_{2}<\dots<j_{s}$. For convenience, define $j_{0}=1$.
Define the matrix $\widehat{g}$ as $P\widehat{g}=g$, where $P$ is the matrix of cyclic permutation with sign
Proposition 1 implies
$\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
if neither $i$ no $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
The moment map equations give the following identity
$x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c$
where $x_{j}^{P}=x_{j-1}$ with $x_{0}=x_{n}$, and equations
(41) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}$
Define
$\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
Theorem 2.
The following holds:
  • $\widehat{\lambda}_{i}\neq 0,\lambda_{i}$, for $i\neq j_{a}$, $a=0,\dots,s$.
  • (42) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})},$
  • (43) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}$
    (44) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}$
Proof.
Write the moment map equation as
$\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
For the characteristic polynomial of the right side, using lemma 1 from the appendix A we obtain
$\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
Taking $z=x_{j_{a}}+c$ we obtain
$\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
which proves the formula (42) for $|\psi_{j_{a}}|^{2}$.
Similarly, writing the moment map equation as
$\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
Finally, the moment map equations (41) give
$\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives (44). ∎
Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer.
Remark 3.
We can write $g$ as the product
(45) $g=\alpha\,P\,Y,$
where $P$ is the cyclic permutation matrix (40).
Let $\beta$ be an element of the stabilizer $x$, i.e. a diagonal unitary matrix.
We use transformations
(46) $g\to\beta g\beta^{-1}$
with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real.
Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
$\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$.
Thus, we can fix the section of the action of the stabilizer group $H\subset SU(n)$:
(47) $g=\alpha\,P\,Y,$
where
$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\}$
(48) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad i\notin\{j_{0},\ldots,j_{s}\}$
(50) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a},$
and $P$ is defined (40).
Note that with this choice, we have
(51) $\mathrm{det}\,\alpha=\mathrm{det}\,P=\mathrm{det}\,Y=1$
and real matrix $Y$ is othogonal
$YY^{t}=I.$
Define $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1},$ then due to (36)
(52) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c).$
Thus, we described coordinates $(y_{j_{a}},\alpha_{j_{a}})$ on $s$-dimention strata: $y_{j_{a}}>c$ and $\{\alpha_{j_{a}}\}_{a=0}^{s}$ with constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$.
@@ 3. Low–dimensional tori
(The fixed point). In the case when all $\psi_{i}=0$ for $i\geq 2$, we have
@@
(53) $g=P.$
In terms of original CMS coordinates we have $q_{j}=\frac{2\pi j}{n}$. The action of the center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}\mod(2\pi)$.
@@
(One-domensional tori). Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ when $i\neq 1,j$.
@@
  • For $y=x_{j-1}-x_{j}$, $y>c$; $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$, and with $\sum_{i=1}^{n}x_{i}=0$:
    (54) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c),$
    where $\delta_{i\geq j}$ is a step function.
  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j},1,\ldots,1\right)$.
(55) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c}$
@@
(56) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix},$
@@
Note that $r^{2}+q^{2}=1$. It is clear that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.
Remark 4.
Theorem 2 can be obtained from theorem 1 directly using the appropriate limit. In particular, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$ ($j_{0}=1$)
(58) $|\psi_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a},$
(59) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{for }i=j_{a},$
where in the LHS $|\psi_{i}|^{2}$ from (17) and in the RHS $|\psi_{j_{a}}|^{2}$ from (42), and
(60) $|\lambda_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a}-1,$

### v2

@@ 3.2. The maximal stratum
where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
@@
(47) $\omega=\exp\left(\frac{2\pi i}{n}k\right){\rm Id},\qquad\text{ where }k=0,\ldots,n-1.$
@@
If $\gamma_{j}$ are eigenvalues of the matrix $g$, such multiplication changes $\gamma_{j}\mapsto\omega\gamma_{j}$. Eigenvalues of $g$ are exponents of coordinates of the CMS model $\gamma_{j}=\exp(iq_{j})$. Thus the multiplication of $g$ by $\omega$ shifts coordinates as
(48) $q_{j}\mapsto q_{j}+\frac{2\pi}{n}k\ (\mathrm{mod}\ 2\pi)\ .$
Thus, we described the maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}\subset S(\mathcal{O})$. It has a natural coordinate system, $\{x_{i}\}$ that lie inside of the shifted principal Weyl chamber $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\{\alpha_{i}\}$, $|\alpha_{i}|=1$ and $\prod_{i=1}^{n}\alpha_{i}=1$.
3.3. Low–dimensional strata $S(\mathcal{O})_{\{j\}}$
To describe all strata, we need the following technical fact.
Proposition 1.
If $\psi_{k}=0$ for some $k=2,\dots,n$, then the only nonzero matrix element of $g$ in the (k1)(k-1)-th row and the $k$-th column is $g_{k-1,k}$ and $|g_{k-1,k}|=1$.
Proof.
The equation (41) for $j=k$ becomes
(49) $(x_{k}-x_{i}+c)g_{ik}=0.$
If $g_{ik}=0$ for all $i$, then $g$ not invertible. Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$. Thus, in the equation above $g_{k-1,k}$ can be non-zero. For $j\neq k-1$ we have $x_{j}-x_{k}+c\neq 0$, and therefore $g_{jk}=0$. This means that the $k$-th column has only one non-zero element $g_{k-1,k}$.
The unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the (k1)(k-1)-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$. ∎
Fix numbers $\{j\}=\{j_{0}=1,j_{1},\dots,j_{s}\}$ such that $1<j_{1}<j_{2}<\dots<j_{s}\leq n$.
Definition 1.
The stratum $S(\mathcal{O})_{\{j\}}\subset S(\mathcal{O})$ is the subspace where $\psi_{k}\neq 0$ when $k\in\{j_{0},j_{1},\dots,j_{s}\}$ and $\psi_{k}=0$ otherwise. Equivalently, $x_{k}-x_{k+1}>c$ if $k\in\{j_{0},j_{1},\dots,j_{s}\}$ and $x_{k}-x_{k+1}=c$ otherwise.
It is clear that $S(\mathcal{O})_{\{j\}}=\pi^{-1}(B(\mathcal{O})_{\{j\}})$.
Define $\widehat{g}=P^{-1}g$, where $P$ is the signed matrix of cyclic permutation
Proposition 1 implies
(51) $\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
if neither $i$ nor $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
The moment map equation (12) give the following identity
(52) $x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c,$
where $x^{P}=P^{-1}xP$ is the diagonal matrix with entries $x_{j}^{P}=x_{j-1}$ and $x_{0}=x_{n}$. For matrix elements of $\widehat{g}$ we obtain equations
(53) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}.$
Define
(54) $\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
Theorem 2.
The following holds:
  • $\widehat{\lambda}_{i}=0$, if $i\neq j_{a}$, $a=0,\dots,s$.
  • (55) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})}.$
  • (56) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}.$
    (57) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}.$
Proof.
Write the moment map equation as
(58) $\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
For the characteristic polynomial of the right side, using lemma 1 from appendix A we obtain
(59) $\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
Evaluation of this identity at $z=x_{j_{a}}+c$ gives
(60) $\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
which proves the formula (55) for $|\psi_{j_{a}}|^{2}$.
Similarly, writing the moment map equation as
(61) $\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
Finally, the moment map equations (53) give
(62) $\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives formula (57). ∎
Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer of $x$.
Corollary 2.
We can choose a representative $(x,g)$ of the conjugacy class $(x,g)]\in\mu^{-1}(\mathcal{O})/G$, where $x$ is diagonal as above and $g$ is the product
(63) $g=\alpha\,P\,Y.$
Here $P$ is the cyclic permutation matrix (50), $Y$ is
$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\},$
(64) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad\,\ i\notin\{j_{0},\ldots,j_{s}\},$
and $\alpha$ is the following diagonal matrix
(66) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a}.$
Indeed, let $\beta$ be an element of the stabilizer of $x$, i.e., a diagonal unitary matrix. We already used transformations
(67) $g\mapsto\beta g\beta^{-1}$
with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real. Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
(68) $\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$. This proves corollary 2.
One can show that
(69) $\mathrm{det}\,Y=1$
and that the real-valued matrix $Y$ is orthogonal
(70) $YY^{t}={\rm Id}.$
Note that because $\det(P)=1$, we have $\det(\alpha)=1$.
Corollary 3.
$S(\mathcal{O})_{\{j\}}\simeq\mathbb{T}^{s}\times\mathbb{R}^{s}$ with $x_{j_{a}}$ being coordinates on $\mathbb{R}^{s}$ and $\alpha_{j_{a}}$ being coordinates on $\mathbb{T}^{s}$.
3.4. The dynamical stratification of the phase space
The results of the previous sections imply that we have a stratification of $S(\mathcal{O})$ given by the preimages of walls of the shifted Weyl chamber:
@@ 3.4. The dynamical stratification of the phase space
The first part in this disjoint union is the fixed point $S(\mathcal{O})_{\{1\}}$ corresponding to $s=0$. Each stratum $S(\mathcal{O})_{\{j_{0},\dots,j_{s}\}}$ has dimension $2s$. The maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ has maximal dimension 2n22n-2.
This is a stratification where strata are parametrized by the partially ordered set $A=\{1\}\cup\dots\cup\{1<j_{1}<\dots<j_{s}\}$ with the partial order $\{k\}\leq\{j\}$ if $\{k\}$ is a subset of $\{j\}$. , see corollary 3, and
(72) $\overline{S(\mathcal{O})_{\{j\}}}=\sqcup_{\{k\}\leq\{j\}}S(\mathcal{O})_{\{k\}}.$
This property can be checked by taking the appropriate limit. For example, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$
(73) $|\psi_{i}|^{2}$ $\to 0,\qquad\qquad\text{for }i\neq j_{a},$
(74) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{ for }i=j_{a}.$
Here on the left hand side $|\psi_{i}|^{2}$ is taken from (26), on the right hand side $|\psi_{j_{a}}|^{2}$ is taken from (55), and
(75) $|\lambda_{i}|^{2}$ $\to 0,\qquad\qquad\text{for }i\neq j_{a}-1,$
3.4.1. Natural coordinates on $S(\mathcal{O})_{\{j\}}$
On the the maximum stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ define variable $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
@@ 3.4.1. Natural coordinates on $S(\mathcal{O})_{\{j\}}$
Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\ln\alpha_{k}$ for $j=2,\ldots,n$. Then,
@@
This gives a convenient set of coordinates on the maximal stratum: real coordinates $(y_{2},\ldots,y_{n})$ with constraints $y_{i}>c$ and coordinates $(\theta_{2},\ldots,\theta_{n})$ on unit tori. In section 4 we will show that $(y,\theta)$ are the action-angle variables.
On $S(\mathcal{O})_{\{j\}}$ define variables $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1}$ then
(79) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c),$
where $\delta_{i\geq j}$ is a step function, $\delta_{i\geq j}=0$ if $i<j$ and $\delta_{i\geq j}=1$ if $i\geq j$.
Set $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ and define $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$. Thus, we described coordinates $(y_{j_{a}},\theta_{j_{a}})$ with $y_{j_{a}}>c$ on $S(\mathcal{O})_{\{j\}}$. We will show that these are action-angle variables on this stratum.
Because $\{y\}$ are coordinates on the base $B(\mathcal{O})_{\{j\}}$ of the Lagrangian fibration (22) of $S(\mathcal{O})_{\{j\}}$, they can be taken as action variables. In the next section we will show that $\{\theta\}$ are the corresponding angle variables.
3.4.2. Two examples of low–dimensional strata
@@
The fixed point $S(\mathcal{O})_{\{1\}}$. In this case all $\psi_{i}=0$ for $i\geq 2$, i.e., $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$, and we have
@@
  • $g=P.$
In terms of original CMS coordinates we have $q_{j}=\frac{2\pi j}{n}$. The center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}({\rm mod}\ 2\pi)$.
@@
Two–dimensional strata $S(\mathcal{O})_{\{1,j\}}$. Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ only when $i\neq 1,j$. The following is a consequence of theorem 2 and corollary 2
@@
  • $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$.
  • Define $y=x_{j-1}-x_{j}$, then $y>c$. Together with $\sum_{i=1}^{n}x_{i}=0$ this gives:
    (80) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c).$
  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j}=\alpha_{1}^{-1},1,\ldots,1\right)$.
(81) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad\ \quad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c},$
@@
(82) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix}\$,
@@
Note that $r^{2}+q^{2}=1$, which implies that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.